Suppose that you invest $1,000,000 into a stock with an expected return of 8% per year. What is the expected number of years for the investment to grow to $4,000,000?
Answer #1


solution: FV FV=PV*(1+r)^n 4000000=1000000*(1.08)^n 1.08^n=4000000/1000000 1.08^n=4 n log 1.08= log 4 n = log 4/log 1.08 n=0.